OpenStudy (anonymous):

Find the difference. nine-tenths minus two-fifths one-fifth one-half seven-tenths one and two-fifths

OpenStudy (anonymous):

9/10 - 2/5 = 2x2= 4 5x2=10 =9/10 - 4/10 = 5/10= 1/2

OpenStudy (anonymous):

Find the difference. nine-tenths minus two-fifths